factorise 4a 2 -9b 2 -2a -3b

factorise 4a2 -9b2 -2a -3b

Grade:10

3 Answers

Sumit Majumdar IIT Delhi
8 years ago
Dear student,
The solution goes as:
4a^2-9b^2-\left ( 2a+3b \right )=\left ( 2a+3b \right )\left ( 2a-3b \right )-\left ( 2a+3b \right )=\left ( 2a+3b \right )\left ( 2a-3b+1 \right )Regards
Rahul Jiji George
8 years ago
sir,
shouldn’t it be (2a + 3b) (2a -3b – 1)
if not why?
 
Sumit Majumdar IIT Delhi
8 years ago
yes u are correct, it was a typing mistake.
?4a^2-9b^2-\left ( 2a+3b \right )=\left ( 2a+3b \right )\left ( 2a-3b \right )-\left ( 2a+3b \right )=\left ( 2a+3b \right )\left ( 2a-3b-1 \right )